# Deep Improvement Supervision

### v2

@@ 4.3 Algorithm
In standard TRM training, the model attempts to satisfy Theorem 4.1 implicitly. It tries to find a latent configuration $\mathbf{z}_{L}$ such that the final margin is positive. This is often unstable because the gradient signal must traverse the entire recursive chain.
DIS solves this by explicitly enforcing a positive Advantage Margin at every recursive step. We achieve this by constructing a sequence of intermediate targets that strictly contract towards the solution. Let $\{\mathbf{y}_{s}^{\dagger}\}_{s=1}^{N_{\text{sup}}}$—constructed so that the expected discrepancy to $\mathbf{y}^{\star}$ strictly decreases with $s$. Let $\mathbf{x}\!\in\!\mathcal{X}$ be the input, $\mathbf{y}^{\star}\!\in\!\mathcal{Y}$ the final target, and $(\mathbf{y},\mathbf{z})$ the TRM state passed across supervision steps. A target generator $\Phi$ produces a sequence of stepwise targets
with $\mathbf{y}^{\dagger}_{N_{\text{sup}}}=\mathbf{y}^{\star}$, such that the distance to the final solution decreases monotonically along the schedule, e.g. where $d$ is a task-appropriate discrepancy (e.g., Hamming distance over tokens). Each supervision step $s\in\{1,\ldots,N_{\text{sup}}\}$ is indexed and receives a puzzle embedding $E(p)$. Our algorithm admits diverse sources of intermediate targets $\Phi$:
  1. 1.
    Programmable edits. Deterministic code-based generator creates path from input to output, for example puzzle solvers that reveal $N$ constraint-consistent move per step, yielding $\mathbf{y}^{\dagger}_{s+1}=\mathrm{Edit}(\mathbf{y}^{\dagger}_{s})$.
  2. 2.
    LLM-generated plans. A teacher LLM proposes a sequence of intermediate solutions or sketches; these are projected onto the task’s discrete output space to form $\{\mathbf{y}^{\dagger}_{s}\}$.
  3. 3.
    Discrete diffusion schedules. Define a corruption process $q_{\beta}(\tilde{\mathbf{y}}\,|\,\mathbf{y}^{\star})$ (e.g., token masking or random replacement with rate $\beta$). Choose a decreasing noise schedule and sample $\mathbf{y}^{\dagger}_{s}\sim q_{\beta_{s}}(\cdot\,|\,\mathbf{y}^{\star})$ so that the targets become progressively less corrupted, approximating a reverse-diffusion path over discrete outputs [^2] We do not position our method as diffusion models. It does not employ diffusion-based schedulers, nor does it explicitly compute concrete scores or distributional ratios (lou2023discrete). We propose a general framework for guiding the reasoning steps..
These constructions guarantee by design (or in expectation) that the improvement direction is explicit. In our paper, we choose the simplest version with stepwise targets via discrete corruption. Let $\mathbf{x}$ be the input and $\mathbf{y}^{\star}$ the ground-truth output. Choose a decreasing noise schedule $0=\beta_{N_{\text{sup}}}<\cdots<\beta_{2}<\beta_{1}\leq 1$ and a token-level corruption kernel $q_{\beta}$ (e.g., masking at rate $\beta$). Define a sequence of intermediate targets
so that $\mathbf{y}^{\dagger}_{N_{\text{sup}}}=\mathbf{y}^{\star}$ and, in expectation, the discrepancy with $\mathbf{y}^{\star}$ (e.g., the Hamming distance) decreases with $s$. We pass $(\mathbf{y}_{s+1},\mathbf{z}_{s+1})$ to the next deep supervision step:
During training, the model receives step-by-step supervision targets and computes losses at each improvement step, allowing it to learn how to sequentially enhance its answers through backpropagation across the reasoning chain.
In this framework, a diffusion corruption process is merely one option for supervising the reasoning process. As demonstrated above, any other sampler for improvement guidance may be used. Crucially, we incorporate the time step t into the model input and observe that an integer-based time index $(0,1\dots,N_{sup})$ yields superior results compared to standard continuous diffusion time conditioning $t\in[0,1])$. Simplification vs. TRM are:
  • Recursion budget: DIS: T=1,n=2T{=}1,n{=}2 vs. TRM: T=3,n=6T{=}3,n{=}6 on ARC—DIS backpropagates through one cycle with two internal latent/answer updates; TRM runs no-grad cycles before a grad cycle. The total step formula is $N_{sup}×[T×(n+1)]$, so TRM does 16[3(6+1)]=33616*[3*(6+1)]=336, while we have $6×(2+1)=\textbf{18}$
  • Supervision: DIS trains each step toward a step-specific target $\mathbf{y}_{s}^{\dagger}$ that provably induces monotone improvement; TRM supervises every step directly on the final label $\mathbf{y}^{\star}$.
  • Halting/ACT: DIS uses a fixed $N_{\text{sup}}{=}6$ with no halting head and no extra forward pass; TRM/HRM use halting (HRM’s ACT requires a second forward pass for the continue loss).
  • Backbone & pipeline: We keep TRM’s attention backbone and augmentation/evaluation pipeline for a fair comparison on ARC, as self-attention generalizes better on $30{\times}30$ puzzles.
Our DIS experiments adopt a minimal-compute design: we retain TRM’s tiny 2-layer attention backbone and ARC protocol but use only six supervised improvement steps with a single external cycle (comprising two internal updates) and no halting head. This streamlined setup isolates the contribution of guided, monotonic improvement.
5 Experiments
In this section, we provide a detailed explanation and results on the complex N-Queens, Sudoku-Extreme and ARC-AGI problems.
Backbone. Our DIS model reuses the tiny single-network TRM backbone but eliminates TRM’s extra recursion and halting heads. We use a 2-layer Transformer block with RMSNorm, SwiGLU MLPs, and rotary position embeddings; weights are shared for both the latent-update and answer-update calls, exactly as in TRM’s attention variant (“TRM-Att”), to isolate the contribution of DIS from capacity and architecture differences (jolicoeur2025less). The task specific hyperparameters as the hidden layers size and reasoning steps are presented below, per task protocol.
5.1 N-Queens
Task Format. The N-Queens problem is a combinatorial reasoning task that involves placing $Q$ queens on a $8\times 8$ chessboard (oarga2025generalizable). The fundamental objective is to arrange the queens such that no two queens threaten each other, which imposes the strict constraint that no two queens can share the same row, column, or diagonal.
Only the classic $8\times 8$ board and no augmentation was used, an example of input and target is represented in 4. We generated 7,200 train and 200 test examples with a sequence length of 64 and a vocabulary size of 3.
Figure 5: Accuracy curves on N-Queens problem.
Results. The purpose of this experiment was to assess the impact of DIS training. As mentioned in §4, we used T = 1 and n = 2 for DIS, while the regular TRM of T = 3 and n = 6 was applied, with a halting head and 16 supervision steps. We achieved the same 0.69 accuracy while using much fewer inference steps. The architectures of the 0.8 mill parameters were used for both methods.
5.2 ARC Evaluation Protocol
Data augmentation. We adopt the augmentation pipeline of TRM to mitigate small-data overfitting: 1000 augmentations per puzzle via color permutations, dihedral-group transforms (90° rotations, flips/reflections) and translations. As in TRM, we also include the 160 ConceptARC tasks as additional training puzzles (moskvichev2023conceptarc). We attach a puzzle-specific embedding token per instance.
5.3 ARC Model Settings
Objective. Each supervision step $s\in\{1,\dots,6\}$ is trained toward a step-specific intermediate target $\mathbf{y}^{\dagger}_{s}$ produced by a discrete corruption schedule of ground truth $\mathbf{y}^{\star}$ with monotonically decreasing noise. We use token-masking/replacement with a linearly decreasing mask rate over the 6 steps so that $\mathbb{E}[d(\mathbf{y}^{\dagger}_{s},\mathbf{y}^{\star})]$ decreases with $s$. The loss is standard token-level cross-entropy on $f_{O}(\mathbf{y})$ against $\mathbf{y}_{s}^{\dagger}$, with linearly increasing step weights $w_{s}$ to emphasize late-step fidelity.
Optimization. We follow TRM’s stable training recipe wherever applicable: Adam-Atan with $\beta_{1}{=}0.9,\ \beta_{2}{=}0.95$, a 2k-step warm-up, and the stable-max cross-entropy variant for stability. For ARC experiments, we use weight decay 0.1 and we did not find EMA important.
We match the hidden size of the TRM D=512D{=}512 and call it the medium model in Table 1. When we use D=256D{=}256 and the single decoder layer model, which results in 0.8 mil. parameters, we call it compact. Also, we match batch sizing; embedding LR warm-up and an elevated embedding LR (as in TRM) are retained.
Deep improvement supervision loop. For each mini-batch we run $N_{\text{sup}}{=}6$ DIS steps. At each step, we execute a single external cycle (since T=1T{=}1) comprising two internal latent/answer updates (n=2n{=}2), backpropagating through the full cycle; we then detach $(\mathbf{y},\mathbf{z})$ before the next step. We do not train a halting/ACT head.
Test-time compute. We run the same $N_{\text{sup}}{=}6$ steps at evaluation. To compare fairly with prior ARC protocols, we keep TRM’s test-time augmentation vote: run the model across 1000 geometric/color augmentations of a puzzle and return the most common prediction.
5.4 ARC Results
Table 1: Model Performance Comparison, pass@2
Method Params ARC-1 ARC-2
Chain-of-thought, pretrained
Deepseek R1 671B 15.8 1.3
Claude 3.7 16K ? 28.6 0.7
o3-mini-high ? 34.5 3.0
Gemini 2.5 Pro 32K ? 37.0 4.9
Grok-4-thinking 1.7T 66.7 16.0
Bespoke (Grok-4) 1.7T 79.6 29.4
Small-sample training
TRM-compact 0.8M 12.0 0.0
DIS-compact (Ours) 0.8M 24.0 0.0
TRM-medium 7M 27.1 0.0
DIS-medium (Ours) 7M 40.0 2.5
TRM 7M 40.4 3.3
DIS 7M 41.3 ?
Shaped credit assignment across supervision steps. In baseline TRM, every step is trained directly against $\mathbf{y}^{\star}$, leaving it to the model to discover a self-improvement curriculum DIS supplies explicit intermediate targets $\{\mathbf{y}^{\dagger}_{s}\}$, aligning the step-$s$ gradients with a concrete improvement objective. This reduces the burden on the latent state $\mathbf{z}$ to implicitly encode a stepwise plan and can accelerate optimization in scarce-data regimes, where TRM has been shown to be the most effective. DIS retains TRM’s minimal two-feature interface $(\mathbf{y},\mathbf{z})$, single tiny network reused for both updates, and the schedule of T1T{-}1 no-grad cycles followed by one grad cycle. It inherits the simplicity advantages of TRM while changing only the supervision signal.
Compute and stability. With a monotone schedule, DIS turns each supervision step into a measurable sub-goal. We preserves TRM’s compute profile per step (one gradient-bearing recursion cycle) and and avoid HRM/TRM-style ACT. If targets are generated offline, the runtime overhead is negligible; if produced online (e.g., by a teacher model), they can be cached or amortized across epochs. For training we used the same 4 GPU H100 setting as TRM, but learning takes $\approx 40$ hours against $72$ in TRM.
6 Discussion

### v3

@@ 4.2 Algorithm
Standard TRM training backpropagates a single terminal supervision signal through the entire recursive chain, so the model must simultaneously (i) discover useful intermediate states and (ii) align each latent update with improvement. DIS isolates the second requirement by supervising the improvement operator itself: each recursion step is trained to map “previous target” $\mathbf{y}^{\dagger}_{s-1}$ to “next target” $\mathbf{y}^{\dagger}_{s}$, directly instantiating the stepwise improvement interpretation from Section 3.
Let $\{\mathbf{y}_{s}^{\dagger}\}_{s=1}^{N_{\text{sup}}}$ be constructed so that the expected discrepancy to $\mathbf{y}^{\star}$ strictly decreases with $s$. Let $\mathbf{x}\!\in\!\mathcal{X}$ be the input, $\mathbf{y}^{\star}\!\in\!\mathcal{Y}$ the final target, and $(\mathbf{y},\mathbf{z})$ the TRM state passed across supervision steps. A target generator $\Phi$ produces a sequence of stepwise targets
with $\mathbf{y}^{\dagger}_{N_{\text{sup}}}=\mathbf{y}^{\star}$, such that the distance to the final solution decreases monotonically along the schedule under a task-appropriate discrepancy $d$ (e.g., Hamming distance over tokens). Each supervision step $s\in\{1,\ldots,N_{\text{sup}}\}$ receives a puzzle embedding $E(p)$.
Our framework admits diverse sources of intermediate targets $\Phi$ (any mechanism that yields monotone improvement targets is valid, since DIS only requires stepwise improvement, not a particular generative process):
  1. 1.
    Programmable edits. A deterministic generator creates a path from input to output, e.g., puzzle solvers that reveal a constraint-consistent move per step, yielding $\mathbf{y}^{\dagger}_{s+1}=\mathrm{Edit}(\mathbf{y}^{\dagger}_{s})$.
  2. 2.
    LLM-generated plans. A teacher LLM proposes intermediate sketches; these are projected onto the task’s discrete output space to form $\{\mathbf{y}^{\dagger}_{s}\}$.
  3. 3.
    Discrete corruption schedule. Define a corruption process $q_{\beta}(\tilde{\mathbf{y}}\,|\,\mathbf{y}^{\star})$ (e.g., token masking or random replacement with rate $\beta$). Choose a decreasing noise schedule and sample $\mathbf{y}^{\dagger}_{s}\sim q_{\beta_{s}}(\cdot\,|\,\mathbf{y}^{\star})$ so that the targets become progressively less corrupted, yielding a simple monotone-improvement curriculum. [^1] We do not position DIS as diffusion modeling: DIS does not require diffusion schedulers, score matching, or explicit density-ratio estimation. A corruption schedule is merely a convenient way to generate monotone intermediate targets in discrete spaces (Lou et al., 2023).
In our experiments, we use the simplest instantiation: stepwise targets via discrete corruption. Let $\mathbf{x}$ be the input and $\mathbf{y}^{\star}$ the ground-truth output. Choose a decreasing noise schedule $0=\beta_{N_{\text{sup}}}<\cdots<\beta_{2}<\beta_{1}\leq 1$ and a token-level corruption kernel $q_{\beta}$ (e.g., masking at rate $\beta$). Define intermediate targets
so that $\mathbf{y}^{\dagger}_{N_{\text{sup}}}=\mathbf{y}^{\star}$ and, in expectation, the discrepancy with $\mathbf{y}^{\star}$ decreases with $s$. Please, see Appendix A for details.
During training, the model runs the TRM recursion for $N_{\text{sup}}$ supervision steps. At each step $s$, it computes the post-reasoning logits $\ell_{c}^{s}$ and applies $\mathrm{CE}(\ell_{c}^{s},\mathbf{y}^{\dagger}_{s})$. By the theory in Section 3, this stepwise supervision trains each latent update to increase the relative log-probability of the next improved target – i.e., to implement a non-trivial policy improvement operator rather than an unconstrained recurrent computation.
5 Experiments
In this section, we provide a detailed explanation and results on the complex N-Queens and ARC-AGI problems.
Backbone. Our DIS model reuses the tiny single-network TRM backbone but eliminates TRM’s extra recursion and halting heads. We use a 2-layer Transformer block with RMSNorm, SwiGLU MLPs, and rotary position embeddings; weights are shared for both the latent-update and answer-update calls, exactly as in TRM’s attention variant (“TRM-Att”), to isolate the contribution of DIS from capacity and architecture differences (Jolicoeur-Martineau, 2025). The task specific hyperparameters as the hidden layers size and reasoning steps are presented below, per task protocol.
From algorithmic perspective, we incorporate the time step $t$ into the model input and observe that an integer-based time index $(0,1\dots,N_{sup})$ yields superior results compared to standard continuous diffusion time conditioning $t\in[0,1])$. The total simplification that we made in comparison to the original TRM are:
  • Recursion budget: DIS: T=1,n=2T{=}1,n{=}2 vs. TRM: T=3,n=6T{=}3,n{=}6 The total step formula is $N_{sup}×[T×(n+1)]$, so TRM does 16[3(6+1)]=33616*[3*(6+1)]=336, while we have $6×(2+1)=\textbf{18}$
  • Halting/ACT: Although TRM/HRM use halting where HRM’s ACT requires a second forward pass for the continue loss, DIS uses a fixed $N_{\text{sup}}{=}6$ with no halting head and no extra forward pass.
In our experiments, we match the hidden size of the TRM D=512D{=}512 and call it the medium model in Table 1. When we use D=256D{=}256 and the single decoder layer model, which results in 0.8 mil. parameters, we call it compact.
5.1 N-Queens
Task Format. The N-Queens problem is a combinatorial reasoning task that involves placing $Q$ queens on a $8\times 8$ chessboard (Oarga and Du, 2025). The fundamental objective is to arrange the queens such that no two queens threaten each other, which imposes the strict constraint that no two queens can share the same row, column, or diagonal.
Only the classic $8\times 8$ board and no augmentation was used, an example of input and target is represented in 1. We generated 7,200 train and 200 test examples with a sequence length of 64 and a vocabulary size of 3.
Figure 2: Accuracy curves on N-Queens problem.
Results. This experiment demonstrates the impact of DIS training. As mentioned in §4, we used T = 1 and n = 2 for DIS, while the regular TRM of T = 3 and n = 6 was applied, with a halting head and 16 supervision steps. We achieved the same 0.69 accuracy while using much fewer inference steps. The architectures of the 0.8 mill parameters were used for both methods.
5.2 ARC Evaluation Protocol
Data augmentation. We adopt the augmentation pipeline of TRM: 1000 augmentations per puzzle via color permutations, dihedral-group transforms (90° rotations, flips/reflections) and translations. As in TRM, we also include the 160 ConceptARC tasks as additional training puzzles (Moskvichev et al., 2023). We attach a puzzle-specific embedding token per instance.
5.3 ARC Model Settings
Objective. Each supervision step $s\in\{1,\dots,6\}$ is trained toward a step-specific intermediate target $\mathbf{y}^{\dagger}_{s}$ produced by a discrete corruption schedule of ground truth $\mathbf{y}^{\star}$ with monotonically decreasing noise. We use token-masking/replacement with a linearly decreasing mask rate over the 6 steps so that $\mathbb{E}[d(\mathbf{y}^{\dagger}_{s},\mathbf{y}^{\star})]$ decreases with $s$. The loss is the standard token-level cross-entropy in $f_{O}(\mathbf{y})$ against $\mathbf{y}_{s}^{\dagger}$, with linearly increasing step weights $w_{s}$ to emphasize late-step fidelity.
Optimization. We follow the training recipe of TRM wherever applicable: Adam-Atan with $\beta_{1}{=}0.9,\ \beta_{2}{=}0.95$, a 2k warm-up and the stable-max cross-entropy variant for stability. For ARC experiments, we use weight decay 0.1 and we did not find EMA important. Also, we match batch sizing; embedding LR warm-up and an elevated embedding LR (as in TRM) are retained.
Deep improvement supervision loop. For each mini-batch we run $N_{\text{sup}}{=}6$ DIS steps. At each step, we execute a single external cycle (since T=1T{=}1) comprising two internal latent/answer updates (n=2n{=}2), backpropagating through the entire cycle; we then detach $(\mathbf{y},\mathbf{z})$ before the next step. We do not train a halting/ACT head.
Test-time compute. We run the same $N_{\text{sup}}{=}6$ steps at evaluation. To compare fairly with previous ARC protocols, we keep TRM’s test-time augmentation vote: run the model over 1000 geometric/color augmentations of a puzzle and return the most common prediction.
5.4 ARC Results
Table 1: Model Performance Comparison, pass@2
Method Params ARC-1 ARC-2
Chain-of-thought, pretrained
Deepseek R1 671B 15.8 1.3
Claude 3.7 16K ? 28.6 0.7
o3-mini-high ? 34.5 3.0
Gemini 2.5 Pro 32K ? 37.0 4.9
Grok-4-thinking 1.7T 66.7 16.0
Bespoke (Grok-4) 1.7T 79.6 29.4
Small-sample training
TRM-compact 0.8M 12.0 0.0
DIS-compact (Ours) 0.8M 24.0 0.0
TRM-medium 7M 27.1 0.0
DIS-medium (Ours) 7M 40.0 2.5
TRM 7M 40.4 3.3
DIS 7M 41.3 6.0
Shaped credit assignment across supervision steps. In baseline TRM, every step is trained directly against $\mathbf{y}^{\star}$, leaving it to the model to discover a self-improvement curriculum DIS supplies explicit intermediate targets $\{\mathbf{y}^{\dagger}_{s}\}$, aligning the step-$s$ gradients with a concrete improvement objective.
This reduces the burden on the latent state $\mathbf{z}$ to implicitly encode a stepwise plan and can accelerate optimization in scarce-data regimes, where TRM has been shown to be the most effective. DIS retains TRM’s minimal two-feature interface $(\mathbf{y},\mathbf{z})$, single tiny network reused for both updates, and the schedule of T1T{-}1 no-grad cycles followed by one grad cycle. It inherits the simplicity advantages of TRM while changing only the supervision signal.
Compute and stability. With a monotone schedule, DIS turns each supervision step into a measurable sub-goal. We preserve the TRM’s compute profile per step (one gradient-bearing recursion cycle) and avoid HRM/TRM-style ACT. If targets are generated offline, the runtime overhead is negligible; if produced online (e.g., by a teacher model), they can be cached or amortized across epochs. For training we used the same 4 GPU H100 setting as TRM, but learning takes $\approx 40$ hours against $72$ in TRM.
6 Conclusion
The key contribution is Deep Improvement Supervision, builds on this insight by introducing a structured, stepwise training regime. DIS provides intermediate targets through a discrete diffusion process, transforming the challenging problem of long-term credit assignment into a more tractable supervised learning task (Ho and Salimans, 2022; Frans et al., 2025).
Our approach not only simplifies training by eliminating the need for learned halting mechanisms but also enhances efficiency, reducing the number of forward passes by 18x with high accuracy.
Appendix A Illustrations
@@
Appendix B Proofs
Proof of Proposition 4.1.
Fix a supervision step $s$ and its state $s_{s}$ (teacher-forced), with previous target $\mathbf{y}^{\dagger}_{s-1}$ and current target $\mathbf{y}^{\dagger}_{s}$. Write the pre/post policies as
$\hat{\pi}_{s}(\cdot\mid s_{s})=\operatorname{softmax}(\ell_{u}^{s}),\qquad\pi_{s}^{+}(\cdot\mid s_{s})=\operatorname{softmax}(\ell_{c}^{s}),$
and define the logit residual $\Delta\ell^{s}(a)=\ell_{c}^{s}[a]-\ell_{u}^{s}[a]$.
Step 1: Relate the margin in logits to the margin in advantages.
Define the log-ratio advantage (Section 3)
$A_{s}(a)\;:=\;\log\pi_{s}^{+}(a\mid s_{s})-\log\hat{\pi}_{s}(a\mid s_{s}).$
Using $\log\operatorname{softmax}(\ell)[a]=\ell[a]-\log\sum_{a^{\prime}}e^{\ell[a^{\prime}]}$, we have
$A_{s}(a)\;=\;\big(\ell_{c}^{s}[a]-\ell_{u}^{s}[a]\big)\;-\;\Big(\log Z_{c}^{s}-\log Z_{u}^{s}\Big)\;=\;\Delta\ell^{s}(a)\;-\;C_{s},$
where $Z_{c}^{s}=\sum_{a^{\prime}}e^{\ell_{c}^{s}[a^{\prime}]}$, $Z_{u}^{s}=\sum_{a^{\prime}}e^{\ell_{u}^{s}[a^{\prime}]}$, and $C_{s}$ does not depend on $a$. Therefore, for any distribution $\mu$ over actions,
$\Delta\ell^{s}(\mathbf{y}^{\dagger}_{s})-\mathbb{E}_{a\sim\mu}\Delta\ell^{s}(a)\;=\;A_{s}(\mathbf{y}^{\dagger}_{s})-\mathbb{E}_{a\sim\mu}A_{s}(a).$
So it suffices to prove the claim for $A_{s}$.
Step 2: DIS makes the post-reasoning policy concentrate on the improved target.
DIS minimizes $\mathrm{CE}(\ell_{c}^{s},\mathbf{y}^{\dagger}_{s})=-\log\pi_{s}^{+}(\mathbf{y}^{\dagger}_{s}\mid s_{s})$ at every step, so along training we drive
$\pi_{s}^{+}(\mathbf{y}^{\dagger}_{s}\mid s_{s})\;\to\;1.$
Equivalently, for any $\varepsilon>0$ there exists a training time after which $\pi_{s}^{+}(\mathbf{y}^{\dagger}_{s}\mid s_{s})\geq 1-\varepsilon$.
Step 3: The interpolated policy $\pi_{s,w}$ also concentrates on the improved target.
Recall $\pi_{s,w}(a\mid s_{s})\propto\hat{\pi}_{s}(a\mid s_{s})^{1-w}\,\pi_{s}^{+}(a\mid s_{s})^{w}$ for any fixed $w\geq 0$. Since softmax policies have full support, $\hat{\pi}_{s}(\mathbf{y}^{\dagger}_{s}\mid s_{s})>0$. Hence the multiplicative form implies: if $\pi_{s}^{+}(\mathbf{y}^{\dagger}_{s}\mid s_{s})\geq 1-\varepsilon$, then also
$\pi_{s,w}(\mathbf{y}^{\dagger}_{s}\mid s_{s})\;\geq\;1-\varepsilon_{w}\qquad\text{with }\varepsilon_{w}\to 0\text{ as }\varepsilon\to 0$
(for fixed $w$, because $\pi_{s,w}$ inherits concentration from $\pi_{s}^{+}$).
Step 4: Concentration implies a positive advantage margin.
Decompose the expectation:
$\mathbb{E}_{a\sim\pi_{s,w}}A_{s}(a)=\pi_{s,w}(\mathbf{y}^{\dagger}_{s})\,A_{s}(\mathbf{y}^{\dagger}_{s})+\sum_{a\neq\mathbf{y}^{\dagger}_{s}}\pi_{s,w}(a)\,A_{s}(a).$
Thus
$A_{s}(\mathbf{y}^{\dagger}_{s})-\mathbb{E}_{a\sim\pi_{s,w}}A_{s}(a)=\sum_{a\neq\mathbf{y}^{\dagger}_{s}}\pi_{s,w}(a)\Big(A_{s}(\mathbf{y}^{\dagger}_{s})-A_{s}(a)\Big).$
Under DIS, $\pi_{s}^{+}(\mathbf{y}^{\dagger}_{s})\to 1$ forces $\log\pi_{s}^{+}(\mathbf{y}^{\dagger}_{s})-\log\pi_{s}^{+}(a)\to+\infty$ for $a\neq\mathbf{y}^{\dagger}_{s}$, while $\hat{\pi}_{s}$ is fixed at that training step; hence $A_{s}(\mathbf{y}^{\dagger}_{s})-A_{s}(a)>0$ for all $a\neq\mathbf{y}^{\dagger}_{s}$ once training is sufficiently advanced. Since $\pi_{s,w}(a)\geq 0$ and $\sum_{a\neq\mathbf{y}^{\dagger}_{s}}\pi_{s,w}(a)=1-\pi_{s,w}(\mathbf{y}^{\dagger}_{s})=\varepsilon_{w}$, we obtain
$A_{s}(\mathbf{y}^{\dagger}_{s})-\mathbb{E}_{a\sim\pi_{s,w}}A_{s}(a)\;>\;0$
as soon as $\varepsilon_{w}$ is small enough and the strict gaps $A_{s}(\mathbf{y}^{\dagger}_{s})-A_{s}(a)$ are positive.
Step 5: Apply at the final target and take expectation.
At the last supervision step $s=N_{\text{sup}}$, $\mathbf{y}^{\dagger}_{s}=\mathbf{y}^{\star}$. Combining Steps 1–4 and taking expectation over data and model stochasticity yields
$\mathbb{E}\Big[\Delta\ell[\mathbf{y}^{\star}]-\mathbb{E}_{a\sim\pi_{w}}\Delta\ell[a]\Big]=\mathbb{E}\Big[A(\mathbf{y}^{\star})-\mathbb{E}_{a\sim\pi_{w}}A(a)\Big]\;>\;0.$
Finally, the assumption $\log\tfrac{P(\mathbf{y}^{\dagger}_{s})}{P(\mathbf{y}^{\dagger}_{s-1})}>0$ ensures the supervision sequence corresponds to genuine stepwise improvements under the task scoring $P$, i.e., the desired direction of advantage is consistent with the target generator rather than arbitrary. ∎
Appendix C Discussion